# LoKO: Low-Rank Kalman Optimizer for Online Fine-Tuning of Large Models

### v1

@@ 1 Introduction
In this paper, we leverage the low-rank decomposition technique from LoRA to reduce the number of trainable parameters in specific layers. We show that the EKF algorithm can be particularly useful for fine-tuning as fine-tuning only involves a small portion of model parameters. We introduce LoKO, a Kalman-based scalable training algorithm, as an alternative to advanced optimizers for online fine-tuning of large models. Particularly, our contributions are:
  • We employ the Low-Rank Adaptation (LoRA) method, introduced by Hu et al. (2021), to significantly reduce the number of trainable parameters, and scale up the Kalman algorithm, enhancing its compatibility for fine-tuning of large models.
  • We adopt a diagonal approximation for the covariance matrix ${\bm{P}}$ to reduce the computational overhead from quadratic to linear and take advantage of the exponential moving average (EMA) for estimating the matrix ${\bm{R}}$ to improve performance without additional computational cost.
  • We conduct various experiments to demonstrate LoKO’s performance in online fine-tuning tasks across computer vision and language modeling domains.
  • To the best of our knowledge, this is the first successful attempt to fine-tune large models, including transformers with millions of parameters and complex architectures, using the Kalman filter algorithm.
@@ 3.2 Low-Rank Adaptation (LoRA)
At the beginning of training, the matrix ${\bm{A}}$ is initialized with a random Gaussian distribution ${\mathcal{N}}(\mathbf{0},\sigma^{2}{\bm{I}})$, and matrix ${\bm{B}}$ is initialized to zero. Using this low-rank decomposition, LoRA reduces the number of trainable parameters from $d\times q$ to $r\times(d+q)$, where $r$ is much smaller than $d$ and $q$. This technique can be applied on certain layers of a neural network model, $h({\bm{x}},{\bm{\theta}})$, resulting in a re-parameterized version, $h_{LoRA}({\bm{x}},\tilde{{\bm{\theta}}})$, with a reduced number of trainable parameters. This reduction enables efficient fine-tuning of large models, making it feasible to apply the Kalman filter.
@@ 5.1 Experiments Setup
We assess the performance of LoKO by implementing it in various well-established computer vision and language models. Our computer vision experiments involve online fine-tuning for image classification on the MNIST dataset using DenseNet-121 with 7 million parameters (Huang et al., 2017), the CIFAR-10/100 dataset with ResNet-18 and ResNet-50 (He et al., 2016), and ViT-B16 (Dosovitskiy, 2020), which contain 12 million, 26 million, and 86 million parameters, respectively. Furthermore, we evaluated ImageNet100 using ViT-L16 (Dosovitskiy, 2020), 305 million parameters. For text classification tasks, we examine the SST-2, CoLA, and MRPC datasets using RoBERTa-base and RoBERTa-large (Liu, 2019), which have 125 million and 355 million parameters, respectively.
We use AdamW (Loshchilov & Hutter, 2017) and AdaGrad (Duchi et al., 2011) as baseline methods since they are widely used in LoRA. We set the batch size to 1 so that the learning algorithms train the models in the context of online classification tasks at each time step. We configure AdamW with a linear learning rate decay schedule (weight_decay=0.0001), starting from 1e-4, which empirically yields the highest accuracy after a preliminary sweep of various learning rates and schedules. In addition, $\beta_{1}$ and $\beta_{2}$ for this optimizer are set to 0.9 and 0.999, respectively. Furthermore, the learning rate of AdaGrad is set to 0.001 with no weight decay. The hyperparameters in our LoKO algorithm are set according to our discussion in 5.3. The diagonally approximated covariance vector $\hat{{\bm{p}}}_{k}$ is initialized with random numbers sampled from a uniform distribution between 0 and 0.2. Furthermore, we use our second method for $\hat{{\bm{R}}}_{k}$ estimation, and the forgetting factor for the EMA is set to $\beta=0.95$. We also adopt the pre-trained models introduced by (Caron et al., 2021) as the backbone for our fine-tuning experiments with ViT. As shown in Table 1, low-rank decomposition is applied on query layers in transformer-based architectures and convolution layers in CNN networks. All experiments are conducted on an A100 GPU platform three times and the average and standard deviation of them are reported.
DenseNet-121 convolution 7.5 M 166 K
ResNet18 convolution 11 M 150 K
ViT-B16 query 86 M 155 K
ResNet50 convolution 24 M 520 K
ViT-L16 query 305 M 400 K
RoBERTa-base query 125 M 666 K
RoBERTa-large query 355 M 1248 K
@@ 5.2 Main results
We use two primary evaluation metrics in our study. First, we calculate the moving average of the loss with a window size of 1000. Second, we measure the average online accuracy, as adopted by Cai et al. (2021), up to the current timestep $k$. The average online accuracy is defined as $\text{acc}(k)=\frac{1}{k}\sum_{i=1}^{k}\mathbbm{1}_{A}({\bm{y}}_{i})$, where $\mathbbm{1}_{\{\cdot\}}(\cdot)$ is an indicator function, and $A$ is the set of Top1 or Top5 prediction. This metric is computed online during the training process and serves to evaluate the algorithm’s capacity to incorporate new knowledge in real-time. We evaluate the convergence speed of LoKO by comparing the number of observations required to achieve equivalent loss or accuracy levels across different algorithms. Figure 12 illustrate LoKO’s performance in online fine-tuning for image and language classification, compared to LoRA with AdamW and AdaGrad, across various well-established models and datasets. The figures illustrate L1 losses for images and cross-entropy losses for texts, along with accuracy during training. These results demonstrate that LoKO consistently matches or surpasses the performance of alternative optimizers. Additionally, Tables 2 and 3 present the average and standard deviation of online accuracy for a single epoch of online fine-tuning. Note that for CoLA and MRPC datasets, we performed multiple epochs due to their smaller dataset sizes. These results demonstrate that across all models and datasets, LoKO consistently outperforms (and in some cases performs almost equivalently) the other methods by achieving the highest average online accuracy with LoKO. The lower standard deviation in LoKO’s results also highlights its more consistent and robust performance compared to other methods. Furthermore, it is important to emphasize that (sub-)optimal values of the trainable parameters can be attained across a range of LoKO hyperparameters, provided they fall within the permissible initialization interval. This contrasts with gradient-based optimizers, whose performance is highly sensitive to the selection of appropriate hyperparameters, particularly the learning rate.
@@
LoKO (ours) ${88.41}_{\pm 0.2}$ $88.21_{\pm 0.1}$ $83.67_{\pm 0.3}$ $82.69_{\pm 0.4}$ $89.17_{\pm 0.3}$ $88.70_{\pm 0.6}$
LoRA/AdamW $\bm{89.55_{\pm 0.1}}$ $\bm{90.66_{\pm 0.2}}$ $\bm{84.41_{\pm 0.1}}$ $\bm{86.19_{\pm 0.9}}$ $\bm{90.95_{\pm 0.3}}$ $\bm{93.04_{\pm 0.4}}$
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionSST-2 / $RoB_{base}$ / $RoB_{base}$ / $RoB_{base}$
Figure 2: Comparison of LoKO (blue) with LoRA/AdamW (red) and LoRA/AdaGrad (green) across various language models and datasets. The top row presents the training loss, while the bottom row illustrates the average online accuracy against the number of data points observed. This figure presents results for RoBERTa-base. For comprehensive results across all models, refer to Appendix D.
@@ 5.3 Ablation of Initialization and Covariance Approximation
To develop a comprehensive understanding of the behavior of the initialization of $\hat{{\bm{p}}}_{0}$, we begin by evaluating its impact on training from scratch using the Kalman filter. The MNIST dataset is employed for experimentation, utilizing the LeNet-5 architecture due to its fast training speed and low computational demands. Our experiments combine two proposed methods for initializing $\hat{{\bm{p}}}_{0}$ with four widely used parameter initialization strategies: Xavier Uniform and Xavier Normal (Glorot & Bengio, 2010), as well as Kaiming Uniform and Kaiming Normal (He et al., 2015). Figure 3 illustrates the average online accuracy of the MNIST dataset after 1000 iterations of training from scratch. As shown, when $\hat{{\bm{p}}}_{0}$ values are set far from the lower and upper bounds, the algorithm diverges, resulting in low prediction accuracy. The results indicate that initializing $\hat{{\bm{p}}}_{0}$ with random values drawn from a uniform distribution (second method) provides a robust and wide range of choices for $\hat{{\bm{p}}}_{0}$ without causing divergence. In Table 4, we further investigate the initialization of $\hat{{\bm{p}}}_{0}$ in the online fine-tuning of various combinations of Dataset-Model using LoKO.
Dataset - Model method 1 method 2
min max min max
MNIST - DenseNet-121 0.0001 ${0.11}_{\pm 0.01}$ 0.0001 ${0.32}_{\pm 0.01}$
CIFAR10 - ViT-B16 0.001 ${105}_{\pm 3}$ 0.001 ${165}_{\pm 3}$
CIFAR100 - ViT-B16 0.001 ${0.2}_{\pm 0.01}$ 0.001 ${0.59}_{\pm 0.01}$
ImageNet100 - ViT-L16 0.001 ${1.0}_{\pm 0.1}$ 0.001 ${2.0}_{\pm 0.1}$
Our proposed methods for approximating ${\bm{R}}_{k}$ are presented in Table 5 alongside other estimation methods from the literature. Additionally, we have included the accuracy results on the MNIST test dataset for comparison.
Ref. Equation Accuracy
MNIST
Puskorius & Feldkamp (1991) $R_{k}=I$ 92.59 %
Murtuza & Chorian (1994)
Singhal & Wu (1988) $R_{k}=Ie^{-k/50}$ 93.65 %
Singhal & Wu (1989)
Chang et al. (2022)
LoKO (method 1) $R_{k}=\beta R_{k-1}+(1-\beta)\hat{R}_{k}$, 93.81 %
LoKO (method 2) $R_{k}=\beta R_{k-1}+(1-\beta)\hat{R}_{k}$, 94.51 %
5.4 Sensitivity analaysis of $\beta$ in EMA
To assess the impact of hyperparameter $\beta$ on LoKO’s performance, we conducted a sensitivity analysis varying $\beta$ values. Through this analysis, we measured the average online accuracy across different $\beta$ settings. The results illustrate that excessively high or low $\beta$ values notably compromise LoKO’s performance. Optimal performance lies within the range of 0.9 to 0.98 for $\beta$. See Appendix C for more details.
@@ 6 Conclusion
In this study, we present the Low-Rank Kalman Optimizer (LoKO), a scalable Kalman-based training algorithm. LoKO offers a comparative alternative to advanced gradient-based optimizers for online fine-tuning scenarios. By leveraging the low-rank adaptation technique, and a diagonal approximation of the covariance matrix and incorporating a novel observation noise estimation technique based on EMA, we reformulate the EKF algorithm accordingly. Empirical evaluations on the benchmark of computer vision and language models, including MNIST, CIFAR-10/100, ImageNet100, SST-2, COLA, and MRPC demonstrate that LoKO consistently achieves superiority over (or at least comparability with) commonly used optimizers, showcasing its potential for efficient online fine-tuning of large models. Although this paper focuses on online fine-tuning in computer vision and language models, further evaluations are necessary across a broader range of domains, including reinforcement learning. Due to space constraints, we leave these explorations for future work.
@@ Appendix B Proofs and Detailed Derivations
Proof of equation 11a: Here, we will show that  equation 11a is equivalent to equation 3a under the assumption of covariance diagonal approximation. For this aim, we need to show that ${\bm{P}}_{k|k-1}{\bm{H}}_{k}^{\top}=\hat{{\bm{p}}}_{k|k-1}\bullet{\bm{H}}_{k}^{\top}$.
@@
For simplicity, let’s drop the subscripts of $\hat{{\bm{p}}}_{k|k-1}$, and ${\bm{H}}_{k}^{\top}$. Now, define a diagonal matrix: ${\bm{P}}=diag([p_{1},p_{2},...,p_{n}])$, where the diagonal elements $[p_{1},p_{2},...,p_{n}]$ are the elements of the vector $\hat{{\bm{p}}}$, and its off-diagonal elements are zeros. Then, the $i^{th}$ row of ${\bm{P}}{\bm{H}}^{\top}$ can be represented as: $\left({\bm{P}}{\bm{H}}^{\top}\right)_{i}=p_{i}\left({\bm{H}}^{\top}\right)_{i}$, where $\left({\bm{H}}^{\top}\right)_{i}$ represents the $i^{th}$ row of matrix ${\bm{H}}^{\top}$. Now, let’s represent the $i^{th}$ row of the transposed Khatri–Rao product of the vector $\hat{{\bm{p}}}$ and matrix ${\bm{H}}^{\top}$: $\left(\hat{{\bm{p}}}\bullet{\bm{H}}^{\top}\right)_{i}=\left(\hat{{\bm{p}}}\right)_{i}\otimes\left({\bm{H}}^{\top}\right)_{i}$. Since $\left(\hat{{\bm{p}}}\right)_{i}=p_{i}$, and $p_{i}$ is a scalar, the equality $p_{i}\otimes\left({\bm{H}}^{\top}\right)_{i}=p_{i}\left({\bm{H}}^{\top}\right)_{i}$ holds true. Consequently, ${\bm{P}}{\bm{H}}^{\top}=\hat{{\bm{p}}}\bullet{\bm{H}}^{\top}$. ∎
Proof of equation 11c: Here also, we will demonstrate that under the assumption of covariance diagonal approximation, the equation 11c is equivalent to equation 3c in vanilla EKF algorithm.

### v2

@@ 1 Introduction
In this paper, we leverage the low-rank decomposition technique from LoRA to reduce the number of trainable parameters in specific layers. We show that the EKF algorithm can be particularly useful for fine-tuning as fine-tuning only involves a small portion of model parameters. We introduce LoKO, a Kalman-based training algorithm, as an alternative to advanced optimizers for online fine-tuning of large models. Particularly, our contributions are:
  • Based on the Low-Rank Adaptation (LoRA) method, introduced by Hu et al. (2021), which significantly reduces the number of trainable parameters, we demonstrated its compatibility with the Kalman filter. Also, we showed that this combination offers faster performance than traditional optimizers in online fine-tuning scenarios.
  • We employ a diagonal approximation for the covariance matrix $\bm{P}$}, a common approach to reduce the computational overhead from quadratic to linear. By integrating this with the exponential moving average (EMA) for estimating the matrix $\bm{R}$} and incorporating it into the LoRA framework, we achieve improved performance without additional computational cost.
  • We conduct various experiments to demonstrate LoKO’s performance in online fine-tuning classification tasks across computer vision and language modeling domains.
  • To the best of our knowledge, this is the first successful attempt to fine-tune large and complex models, including transformers with millions of parameters, using the Kalman filter algorithm.
@@ 3.2 Low-Rank Adaptation (LoRA)
[…]
@@ 5.1 Experiments Setup
We assess the performance of LoKO by implementing it in various well-established computer vision and language models. Our computer vision experiments involve online fine-tuning for image classification on the MNIST dataset using DenseNet-121 with 7 million parameters (Huang et al., 2017), the CIFAR-10/100 dataset with ResNet-18 and ResNet-50 (He et al., 2016), and ViT-B16 (Dosovitskiy, 2020), which contain 12 million, 26 million, and 86 million parameters, respectively. Furthermore, we evaluated ImageNet100 using ViT-L16 (Dosovitskiy, 2020), 305 million parameters. For text classification tasks, we examine the SST-2, CoLA, and MRPC datasets using RoBERTa-base and RoBERTa-large (Liu, 2019), which have 125 million and 355 million parameters, respectively. We employ the following pre-trained models as backbones for our fine-tuning experiments:
•  For MNIST/DenseNet-121, we used a backbone pre-trained on ImageNet via PyTorch Image Models (timm) (Wightman et al., 2023).
•  For CIFAR-10/ResNet18 and CIFAR-100/ResNet50, we employed pre-trained models from (He et al., 2016), which were trained on ImageNet as the backbone.
•  For CIFAR-10/ViT-B16, CIFAR-100/ViT-B16, and ImageNet100/ViT-L16, we used DINOv2 (Oquab et al., 2023), which was pre-trained on ImageNet.
•  For the language tasks we employed RoBERTa pre-trained model from (Liu, 2019).
We use AdamW (Loshchilov & Hutter, 2017) and AdaGrad (Duchi et al., 2011) as baseline methods since they are widely used in LoRA. We set the batch size to 1 so that the learning algorithms train the models in the context of online classification tasks at each time step. We configure AdamW with a linear learning rate decay schedule (weight_decay=0.0001), starting from 1e-4, which empirically yields the highest accuracy after a preliminary sweep of various learning rates and schedules. In addition, $\beta_{1}$ and $\beta_{2}$ for this optimizer are set to 0.9 and 0.999, respectively. Furthermore, the learning rate of AdaGrad is set to 0.001 with no weight decay. The hyperparameters in our LoKO algorithm are set according to our discussion in 5.3. The diagonally approximated covariance vector $\hat{{\bm{p}}}_{k}$ is initialized with random numbers sampled from a uniform distribution between 0 and 0.2. Furthermore, we use our second method for $\hat{{\bm{R}}}_{k}$ estimation, and the forgetting factor for the EMA is set to $\beta=0.95$. We also adopt the pre-trained models introduced by (Caron et al., 2021) as the backbone for our fine-tuning experiments with ViT. The low-rank decomposition is applied on query layers in transformer-based architectures and convolution layers in CNN networks. All experiments are conducted on an A100 GPU platform three times and the average and standard deviation of them are reported.
@@ 5.2 Main results
We use two primary evaluation metrics in our study. First, we calculate the moving average of the loss with a window size of 1000. Second, we measure the average online accuracy, as adopted by Cai et al. (2021), up to the current timestep $k$. The average online accuracy is defined as $\text{acc}(k)=\frac{1}{k}\sum_{i=1}^{k}\mathbbm{1}_{A}({\bm{y}}_{i})$, where $\mathbbm{1}_{\{\cdot\}}(\cdot)$ is an indicator function, and $A$ is the set of Top1 or Top5 prediction. This metric is computed online during the training process and serves to evaluate the algorithm’s capacity to incorporate new knowledge in real-time. We evaluate the convergence speed of LoKO by comparing the number of observations required to achieve equivalent loss or accuracy levels across different algorithms. Figure 12 illustrate LoKO’s performance in online fine-tuning for image and language classification, compared to LoRA with AdamW and AdaGrad, across various well-established models and datasets. The figures illustrate L1 losses for images and cross-entropy losses for texts, along with accuracy during training. These results demonstrate that LoKO consistently matches or surpasses the performance of alternative optimizers. Additionally, Tables 1 and 2 present the average and standard deviation of online accuracy for a single epoch of online fine-tuning. Note that for COLA and MRPC datasets, we performed multiple epochs due to their smaller dataset sizes. These results demonstrate that across all models and datasets, LoKO consistently outperforms (and in some cases performs almost equivalently) the other methods by achieving the highest average online accuracy with LoKO. The lower standard deviation in LoKO’s results also highlights its more consistent and robust performance compared to other methods. Furthermore, it is important to emphasize that (sub-)optimal values of the trainable parameters can be attained across a range of LoKO hyperparameters, provided they fall within the permissible initialization interval. This contrasts with gradient-based optimizers, whose performance is highly sensitive to the selection of appropriate hyperparameters, particularly the learning rate.
Finally, as can be seen in Table 2, LoKO demonstrates slightly lower performance than AdamW on certain language tasks. A justification for this observation is the fact that the AdamW optimizer benefits from the decoupled weight decay, which has contributed to its strong performance in language modeling tasks (Xie & Li, 2024). On the other hand, the gradient-free Kalman algorithm used in LoKO does not incorporate a similar regularization mechanism, which may explain the slightly lower performance in certain cases compared to AdamW.
At the end, we conducted supplementary experiments by applying the Kalman filter on the DoRA variant of LoRA. These experiments were performed on transformer-based models, and the results were compared against those obtained using AdamW and AdaGrad. The corresponding outcomes are presented in Figure 3 and 4 in Appendix D as well as Table 1 and 2. As shown, the results are promising for the Kalman algorithm, demonstrating its compatibility and effectiveness when integrated with this Weight-Decomposed Low-Rank Adaptation technique (Liu et al., 2024).
@@
DoRA/Kalman $99.84_{\pm 0.03}$ $\bm{92.17_{\pm 0.58}}$ $\bm{97.96_{\pm 0.05}}$
DoRA/AdamW $98.82_{\pm 0.02}$ $89.43_{\pm 0.18}$ $94.50_{\pm 0.08}$
DoRA/AdaGrad $99.01_{\pm 0.04}$ $85.39_{\pm 1.11}$ $92.88_{\pm 0.06}$
@@
LoKO (ours) ${88.41}_{\pm 0.2}$ $88.21_{\pm 0.1}$ $83.67_{\pm 0.3}$ $82.69_{\pm 0.4}$ $9.17_{\pm 0.3} 88$$8.70_{\pm 0.6}$
LoRA/AdamW $89.55_{\pm 0.1}$ $\bm{90.66_{\pm 0.2}}$ $\bm{84.41_{\pm 0.1}}$ $\bm{86.19_{\pm 0.9}}$ $90.95_{\pm 0.3}$} $93.04_{\pm 0.4}$}
@@
DoRA/Kalman $8.32_{\pm 0.2} 88$$8.76_{\pm 0.2}$ $84.05_{\pm 0.0}$ $83.33_{\pm 0.1}$ $89.64_{\pm 0.34}$ $89.49_{\pm 0.27}$
DoRA/AdamW $\bm{89.64_{\pm 0.1}}$ $90.50_{\pm 0.6}$} $84.37_{\pm 0.1}$} $85.33_{\pm 1.5}$} $\bm{91.48_{\pm 0.2}}$ $\bm{93.57_{\pm 0.1}}$
DoRA/AdaGrad $87.99_{\pm 0.1}$ $86.75_{\pm 0.7}$ $78.29_{\pm 0.1}$ $79.51_{\pm 0.7}$ $82.02_{\pm 0.4}$ $83.87_{\pm 0.0}$
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionSST-2 / $RoB_{base}$COLA / $RoB_{base}$MRPC / $RoB_{base}$SST-2 / $RoB_{large}$COLA / $RoB_{large}$MRPC / $RoB_{large}$Refer to caption
@@ 5.3 Ablation of Initialization and Covariance Approximation
Initialization of $\hat{{\bm{p}}}_{0}$:
Initializing $\hat{{\bm{p}}}_{0}$ requires prior knowledge of the network weights and their associated uncertainties. In the absence of such knowledge, almost any initial values for $\hat{{\bm{p}}}_{0}$ can be chosen, provided they are not too close to zero, and in some cases, they are not too far from an upper bound. To develop a comprehensive understanding of the behavior of the initialization of $\hat{{\bm{p}}}_{0}$, we begin by evaluating its impact on training from scratch using the Kalman filter. The MNIST dataset is employed for experimentation, utilizing the LeNet-5 architecture due to its fast training speed and low computational demands. Our experiments combine two proposed methods for initializing $\hat{{\bm{p}}}_{0}$ with four widely used parameter initialization strategies: Xavier Uniform and Xavier Normal (Glorot & Bengio, 2010), as well as Kaiming Uniform and Kaiming Normal (He et al., 2015). Figure 6 in Appendix D illustrates the average online accuracy of the MNIST dataset after 1000 iterations of training from scratch. As shown, when $\hat{{\bm{p}}}_{0}$ values are set far from the lower and upper bounds, the algorithm diverges, resulting in low prediction accuracy. The results indicate that initializing $\hat{{\bm{p}}}_{0}$ with random values drawn from a uniform distribution (second method) provides a robust and wide range of choices for $\hat{{\bm{p}}}_{0}$ without causing divergence. Our experiments demonstrate that as long as the initialization falls within an acceptable range, the filter will eventually converge to the optimal values. Notably, the optimal values are not highly sensitive to the initial covariance matrix. To obtain the upper bounds for our case studies, we conducted a series of experiments across a wide range of initial values and tracked the average online accuracy. If the accuracy doesn’t reach a specific threshold within a predetermined number of iterations, we consider the test to have diverged. The boundary values derived from this analysis have been reported in Table 4 in Appendix D, specifically for case studies where an upper bound for the initial values of the covariance matrix is defined. To the best of our knowledge, no general criterion exists for this selection, and only a few studies have addressed the issue of initialization such as (Heimes, 1998; Rivals & Personnaz, 1998).
Approximation of $\bm{R}$}:
Our proposed methods for approximating ${\bm{R}}_{k}$ are presented in Table 5 in Appendix D alongside other estimation methods from the literature. Additionally, we have included the accuracy results on the MNIST test dataset for comparison.
@@ 6 Conclusion
In this study, we present the Low-Rank Kalman Optimizer (LoKO), a Kalman-based training algorithm. LoKO offers a comparative alternative to advanced gradient-based optimizers for online fine-tuning scenarios. By leveraging the low-rank adaptation technique, and a diagonal approximation of the covariance matrix and incorporating a novel observation noise estimation technique based on EMA, we reformulate the EKF algorithm accordingly. Empirical evaluations on the benchmark of computer vision and language models, including MNIST, CIFAR-10/100, ImageNet100, SST-2, COLA, and MRPC demonstrate that LoKO consistently achieves superiority over (or at least comparability with) commonly used optimizers, showcasing its potential for efficient online fine-tuning of large models. Although this paper focuses on online fine-tuning in computer vision and language models for classification tasks, further evaluations are necessary across a broader range of domains, such as reinforcement learning. We leave these explorations for future work.
@@ B.1 Proof for Proposition 1
Proof of equation 11a
: Here, we will show that  equation 11a is equivalent to equation 3a under the assumption of covariance diagonal approximation. For this aim, we need to show that ${\bm{P}}_{k|k-1}{\bm{H}}_{k}^{\top}=\hat{{\bm{p}}}_{k|k-1}\bullet{\bm{H}}_{k}^{\top}$.
@@
For simplicity, let’s drop the subscripts of $\hat{{\bm{p}}}_{k|k-1}$, and ${\bm{H}}_{k}^{\top}$. Now, define a diagonal matrix: ${\bm{P}}=diag([p_{1},p_{2},...,p_{n}])$, where the diagonal elements $p_{1},p_{2},...,p_{n}]$ are the elements of the vector $\hat{{\bm{p}}}$, and its off-diagonal elements are zeros. Then, the $i^{th}$ row of ${\bm{P}}{\bm{H}}^{\top}$ can be represented as: $\left({\bm{P}}{\bm{H}}^{\top}\right)_{i}=p_{i}\left({\bm{H}}^{\top}\right)_{i}$, where $\left({\bm{H}}^{\top}\right)_{i}$ represents the $i^{th}$ row of matrix $\bm{H}$$^{\top}$. Now, let’s represent the $i^{th}$ row of the transposed Khatri–Rao product of the vector $\hat{{\bm{p}}}$ and matrix $\bm{H}$$^{\top}$: $\left(\hat{{\bm{p}}}\bullet{\bm{H}}^{\top}\right)_{i}=\left(\hat{{\bm{p}}}\right)_{i}\otimes\left({\bm{H}}^{\top}\right)_{i}$. Since $\left(\hat{{\bm{p}}}\right)_{i}=p_{i}$, and $p_{i}$ is a scalar, the equality $p_{i}\otimes\left({\bm{H}}^{\top}\right)_{i}=p_{i}\left({\bm{H}}^{\top}\right)_{i}$ holds true. Consequently, ${\bm{P}}{\bm{H}}^{\top}=\hat{{\bm{p}}}\bullet{\bm{H}}^{\top}$. ∎
Proof of equation 11c
: Here also, we will demonstrate that under the assumption of covariance diagonal approximation, the equation 11c is equivalent to equation 3c in vanilla EKF algorithm.